OpenStudy (anonymous):

What is the solution to the equation? 2(x – 7) = 10x + 18 A. x = 4 B. x = 15/8 C. x = -4 D. x = -15/8

OpenStudy (anonymous):

@Ilovecake

OpenStudy (anonymous):

I realy need this done plz ;_;

OpenStudy (ilovecake):

2x - 14 = 10x + 18 2x - 10x = 14 + 18 -8x = 32 x = 32/-8 x = -4

OpenStudy (anonymous):

Last one What is the solution to the inequality? 4x + 7 > –13 A. x < –5 B. x > –5 C. x < 5 D. x > 5

OpenStudy (ilovecake):

Can you help me with GUM after?

OpenStudy (anonymous):

I'll try

OpenStudy (ilovecake):

It is B

OpenStudy (anonymous):

After this I have to eat I'll be back in 30 just tag when needed

OpenStudy (ilovecake):

okay

OpenStudy (anonymous):

80%1!!

OpenStudy (ilovecake):

Good job!

OpenStudy (ilovecake):

I got 80% too